OZONE RECOVERY IN THE PRESENCE OF O_{2}(a^{1}Δ) FOR ATMOSPHERIC STUDIES

\underline{MICHAEL C. HEAVEN}, Department of Chemistry, Emory University, Atlanta, GA 30322; VA-LERIY N. AZYAZOV, P.N. Lebedev Physical Institute, Samara Branch, Samara, Russia, 443029.

Rapid quenching of O_{2}(a^{1}Δ) in O(^{3}P)/O_{2}/O_{3} mixtures has been observed using 248 nm laser photolysis of ozone to produce oxygen atoms and O_{2}(a^{1}Δ) molecules. Collisional removal of O_{2}(a^{1}Δ) was found to be correlated with the presence of O atoms. The kinetics indicate that vibrationally excited ozone is an important O_{2}(a^{1}Δ) quenching agent in O(^{3}P)/O_{2}/O_{3} systems. The impact of this quenching channel on the O_{2}(a^{1}Δ) and O_{3} levels in the atmosphere will be discussed.